OpenStudy (anonymous):

the integral of 1/(x^2+9)

OpenStudy (anonymous):

I cannot use ln, no?

OpenStudy (solomonzelman):

You can do a trig substitution: \(x=3\tan\theta\) and by this, you get: \(dx=3\sec^2\theta ~ d\theta\)

OpenStudy (solomonzelman):

oh oppsies

OpenStudy (solomonzelman):

No ln's

OpenStudy (anonymous):

oh cool. thanks. so it doesn't have to be under a radical for me to use that trig sub?

OpenStudy (solomonzelman):

You can use trig sub whenever, well, if it doesn't draw you into the world of a very messy notebook

OpenStudy (solomonzelman):

Yes, not necessarily in the root or anything of this sort...

OpenStudy (anonymous):

ok thanks!! that makes it easy

OpenStudy (solomonzelman):

Just that trig substitution is a general get-away from thing such as \(\large\color{black}{ \displaystyle \int \frac{1}{\sqrt{x^2+a^2}} dx }\)

OpenStudy (solomonzelman):

but, that certainly doesn't mean you can't use it in such a case.... you can "relize the derivative" or you can do the integration with trig sub... either way

OpenStudy (solomonzelman):

realize (not relize)

OpenStudy (anonymous):

i see! thanks so much!

OpenStudy (solomonzelman):

Yes, no prob, just one quick post:

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \int \frac{1}{x^2+a^2} dx }\) \(\large\color{black}{ \displaystyle x=a\tan\theta }\) \(\large\color{black}{ \displaystyle dx=a\sec^2\theta{~~}d\theta }\) \(\large\color{black}{ \displaystyle \int \frac{1}{\left(a\tan\theta\right)^2+a^2} \left( a\sec^2\theta{~~}d\theta \right) }\) \(\large\color{black}{ \displaystyle \int \frac{a\sec^2\theta}{a^2\tan^2\theta+a^2} {~~}d\theta }\) \(\large\color{black}{ \displaystyle \int \frac{a\sec^2\theta}{a^2\left(\tan^2\theta+1\right)} {~~}d\theta }\) \(\large\color{black}{ \displaystyle \int \frac{a\sec^2\theta}{a^2\left(\sec^2\theta\right)} {~~}d\theta }\) \(\large\color{black}{ \displaystyle \int \frac{1}{a} {~~}d\theta =\frac{1}{a}\theta }\) \(\large\color{black}{ \displaystyle x=a\tan\theta }\) ---> \(\large\color{black}{ \displaystyle x/a=\tan\theta }\) ---> \(\large\color{black}{ \displaystyle \theta=\tan^{-1}\left(\frac{x}{a}\right) }\) \(\large\color{black}{ \displaystyle \int \frac{1}{a} {~~}d\theta =\frac{1}{a}\theta=\frac{1}{a}\tan^{-1}\left(\frac{x}{a}\right) }\)

OpenStudy (solomonzelman):

thus, \(\large\color{black}{ \displaystyle \int \frac{1}{x^2+a^2} dx =\frac{1}{a}\tan^{-1}\left(\frac{x}{a}\right) }\)

OpenStudy (solomonzelman):

well, +C, but that doesn't matter for what I showed.

OpenStudy (solomonzelman):

it took more time for me because I tried to make sure no typos are there... have fun:)

OpenStudy (anonymous):

thanks! i am working it out right now i will compare it to yours when I'm done!

OpenStudy (solomonzelman):

Ok:)

OpenStudy (solomonzelman):

yw